OpenStudy (anonymous):

Find all solutions to the equation. sin2x+sinx=0

Nnesha (nnesha):

is it sin^2x ?

OpenStudy (anonymous):

\[\sin^2x+sinx=0\]

Nnesha (nnesha):

alright that's the quadratic equation factor t

Nnesha (nnesha):

what is the common factor ?

OpenStudy (anonymous):

sin?

Nnesha (nnesha):

right take it out

Nnesha (nnesha):

it's sin(x) so when we take out sinx from sinx what should we get ?

Nnesha (nnesha):

x represent theta here

OpenStudy (anonymous):

\[sinx^2(2)\]

Nnesha (nnesha):

hmm \[\rm \sin^2x +sinx=0\] can be written as \[\large\rm \sin(x) *\sin(x)+\sin(x)=0\] now take out sinx from both terms

Nnesha (nnesha):

sin(x) is common factor

OpenStudy (anonymous):

There is no way to take it out if they are all sin(x)

OpenStudy (anonymous):

Or does this mean the solution is all real numbers?

Nnesha (nnesha):

hmm lets make it simple sin(x) = y \[\huge\rm y^2+y=0\] take out the common factor

Nnesha (nnesha):

let sin(x) = y^^^

OpenStudy (anonymous):

...lol... am I trying to get y=

OpenStudy (anonymous):

Im sorry

Nnesha (nnesha):

we have to find factor of that quadratic equation

Nnesha (nnesha):

what's the common factor ? y^2+y=0 ?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

y*y+y=0

Nnesha (nnesha):

right now take out y from both terms from y^2 and from y what would you get in other words when we take out common factor we basically divide both terms by common factor

OpenStudy (anonymous):

y(y+1)=0

Nnesha (nnesha):

correct now set it equal to zero \[\rm y=0 ~~~~~~~~~~~~y+1 =0\] now we can replace it back with sin(x) \[\sin(x)=0~~~~~~~~~~~~~~~~\sin(x)+1=0\]

OpenStudy (anonymous):

Okay I see what you mean

Nnesha (nnesha):

make sense ? sin(x) is one term not just sin \[\sin x \rightarrow \sin \theta \]

OpenStudy (anonymous):

Understood

Nnesha (nnesha):

alright sin(x)=0 and sin(x)+1 = 0 ^^^ solve for sin(x)

OpenStudy (anonymous):

sin(x)=0,-1

Nnesha (nnesha):

correct and then unit circle!

OpenStudy (anonymous):

180 and 270?

Nnesha (nnesha):

find `all` solutions

OpenStudy (anonymous):

um

OpenStudy (anonymous):

\[180 \le x \le 270\]

Nnesha (nnesha):

180 and 270 are correct but there is another angle where sin(x) =0

OpenStudy (anonymous):

But that goes on forever

Nnesha (nnesha):

ye true :P

Nnesha (nnesha):

\[180 \le x \le 270\] hmm this make sense

OpenStudy (anonymous):

would you rather me say pi and 2pi

Nnesha (nnesha):

well that depends on the statement does it say in radins or degrees? some times the just give in interval ( pi </ x </2pi)

Nnesha (nnesha):

an*

OpenStudy (anonymous):

Find all solutions to the equation. \[\sin^2x+sinx=0\] Thats exactly what it gives me

Nnesha (nnesha):

\[180 \le x \le 270\] solutions between180 to 270 but look at the 0/260 degree

Nnesha (nnesha):

sin(x)=0 at 0 degree and we need `all` solutions i'm not sure what they want like radian or degree i'll go with degree..

Nnesha (nnesha):

gtg and hmm im not sure

Nnesha (nnesha):

@freckles ,-,

OpenStudy (freckles):

what part are you stuck on @kjones0331 ?

OpenStudy (anonymous):

What solution to put as the answer because this is a free response question

OpenStudy (anonymous):

It says all solutions, be we came up with infinite

OpenStudy (freckles):

solve \[\sin^2(x)+\sin(x)=0 \text{ where } x \in [180,270] ?\] this is the question?

OpenStudy (freckles):

First you need to factor sin^2(x)+sin(x)

OpenStudy (freckles):

oh you need all the solutions?

OpenStudy (freckles):

it just says solve sin^2(x)+sin(x)=0?

OpenStudy (anonymous):

wait..

OpenStudy (freckles):

ok well you still need to factor sin^2(x)+sin(x)

OpenStudy (anonymous):

Could \[x \epsilon, [180,270]\] be the answers

OpenStudy (freckles):

ok so have you factors sin^2(x)+sin(x) yet?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Up top

OpenStudy (freckles):

ok set both factors equal to 0

OpenStudy (freckles):

and just use unit circle

OpenStudy (anonymous):

We did, we got \[sinx=0\] and \[sinx=-1\]

OpenStudy (anonymous):

Which on the unit circle could be 180 or 270

OpenStudy (freckles):

ok sin^2(x)+sin(x)=0 sin(x)(sin(x)+1)=0 sin(x)=0 or sin(x)=-1 this should be easy look at the y-coordinates on the unit circle where do you see y is 0 and y is -1 where the angles are between 180 and 270 (inclusive)

OpenStudy (freckles):

and yes exactly 180 or 270 are the solutions

OpenStudy (freckles):

sin(x)=0 when x is 180 deg sin(x)=-1 when x is 270 deg

OpenStudy (anonymous):

How would I write that as an answer though?

OpenStudy (freckles):

sin^2(x)+sin(x)=0 where 180<=x<=270 has solutions x=180 or 270

OpenStudy (freckles):

that is it we already solved it

OpenStudy (anonymous):

So we should leave it in deg? Basically writing:\[180 \le x \le 270\]

OpenStudy (freckles):

you gave the interval in degree the answer should be in degrees if the interval was in radians then you would answer in radians

OpenStudy (anonymous):

But there was no deg or radians in the original equation, that was what Nnesha and I were stuck on. We have the answer, just dont know how to write it.

OpenStudy (freckles):

I thought you said the question was solve sin^2(x)+sin(x)=0 where 180<=x<=270?

OpenStudy (freckles):

is that not the question?

OpenStudy (anonymous):

Im sorry for the misunderstanding.. the original problem is what I typed as the question

OpenStudy (freckles):

so the question is just sin^2(x)+sin(x)=0 and there was no restrictions on x?

OpenStudy (anonymous):

No

OpenStudy (freckles):

what?

OpenStudy (anonymous):

Im sorry, no there are no restrictions

OpenStudy (freckles):

that is what I said and you said no though

OpenStudy (freckles):

\[\sin^2(x)+\sin(x)=0 \\ \sin(x)[\sin(x)+1]=0 \\ \sin(x)=0 \text{ or } \sin(x)=-1 \] ok I would just put the answer in radians first find all the solutions in the first rotation [0,2pi]

OpenStudy (freckles):

and then we will just add 2pi*k where k is an integer to get all the solutions

OpenStudy (freckles):

you actually already found when sin(x)=-1 on [0,2pi] which is when x=3pi/2 so all the solutions to sin(x)=-1 would be x=3pi/2+2pi*k now sin(x)=0 on [0,2pi] has solutions x=0,pi,2pi if you notice the pattern there then you know that all the solutions to sin(x)=0 would be x=k*pi

OpenStudy (freckles):

where k is an integer

OpenStudy (anonymous):

So I would answer the question with x=kpi?

OpenStudy (freckles):

that is one solution you have to also state the other

OpenStudy (anonymous):

I am getting 3pi/2 ... 7pi/2 ... 11pi/2 ... and so on

OpenStudy (freckles):

why not just say for the sin(x)=-1 we have x=3pi/2+2pi*k ?

OpenStudy (anonymous):

Then 3pi/2 wouldnt fit

OpenStudy (freckles):

why not?

OpenStudy (anonymous):

But I could type that seperately

OpenStudy (freckles):

k=0 is an integer

OpenStudy (anonymous):

the 2pi

OpenStudy (freckles):

I don't know what to tell you sin(x)=0 has solutions x=k*pi sin(x)=-1 has solutions x=3pi/2+2pi*k where k is an integer and 3pi/2 does fit in that last one because k=0 is an integer

OpenStudy (freckles):

3pi/2+2pi*(0) 3pi/2+0 3pi/2 <---see 3pi/2+2pi*k does give us 3pi/2

OpenStudy (anonymous):

Oh okay, sorry

OpenStudy (freckles):

the set of integers are {...,-4,-3,-2,-1,0,1,2,3,4,....}

OpenStudy (anonymous):

Thank you so much